OpenStudy (openstudymathhelp):

ANTIDERIVATIVES Consider the function f(x)=10cosx−8sinx. Let F(x) be the antiderivative of f(x) with F(0)=8.

OpenStudy (openstudymathhelp):

The part that is throwing me off, is the F(0)=8

OpenStudy (retireed):

That is an initial condition so that you can solve what the + C is equal to.

OpenStudy (openstudymathhelp):

So do I find the antiderivatve of 10cos-8sinx, and have that equal to 8?

OpenStudy (retireed):

Kind of, I think you need to plug 0 into the antiderivative with + C and set that equal to 8. I think what you said was correct, I was just over thinking it.

OpenStudy (openstudymathhelp):

i\[10\sin(0)+8\cos(0)+C=8\]

OpenStudy (openstudymathhelp):

ah ok I got it, C would then be equal to zero

OpenStudy (retireed):

YES very good.

OpenStudy (retireed):

I hope I remember this correctly.

OpenStudy (openstudymathhelp):

That sounds familiar to what I learned in class today. I think I was mistaking it as a separate function because I was getting a completely different answer. Thanks!

OpenStudy (retireed):

yw